OpenStudy (anonymous):

Method of Completing the square. Help how do I complete the square for x^2 - 10x = 15

OpenStudy (asnaseer):

Note that if you expand \((x-a)^2\) you get:\[(x-a)^2=x^2-2ax+a^2\]We can rearrange this as:\[x^2-2ax=(x-a)^2-a^2\]See if you can apply this to your problem